OpenStudy (anonymous):

If g(x) = x^2 - 2, find g(6). A -2 B 34 C 10 D 36

OpenStudy (anonymous):

B

OpenStudy (anonymous):

g(6) = 6^2 -2 = 36-2 = 34 which is B